OpenStudy (anonymous):

determine if the function has an inverse that is a function. justify your answer. if it does, find the inverse function: f(x)=(x+3)/(x-2)

OpenStudy (anonymous):

ok this takes a bit of algebra

OpenStudy (anonymous):

rewrite as \[y=\frac{x+3}{x-2}\] and then either solve for x or switch x and y and solve for y. i will do it the second way

OpenStudy (anonymous):

\[x=\frac{y+3}{y-2}\] multiply to get rid of the denominator \[x(y-2)=y+3\] multiply out on the left \[xy-2x=y+3\]

OpenStudy (anonymous):

put all the y's on one side \[xy-y=2x+3\] factor out the y \[y(x-1)=2x+3\] divide to get y by itself \[y=\frac{2x+3}{x-1}\]

OpenStudy (anonymous):

and thats the inverse right?

OpenStudy (anonymous):

that is it!

OpenStudy (anonymous):

thanks!:)

OpenStudy (anonymous):

i mean it is it when you write \[f^{-1}=\frac{2x+3}{x-1}\]